solve 3/x^2+x-2 + 3/x-1 = 1/x+2
Answer by mananth(16946)

3/x^2+x-2 + 3/x-1 = 1/x+2
.
3/(x+2)(x-1) +3/x-1 -1/ x+2=0
.
LCM (x+2)(x-1)
3+3(x+2)-1(x-1) / (x+2)(x-1)=0
..
3+3x+6-x+1 / (x+2)(x-1)=0
2x+10 /(x+2)(x-1)=0
x+5 /(x+2)(x-1)=0
x+5=0
x=-5